# Weyl’s Law for compact rank one symmetric spaces

### v1

@@ 7. Error is sharp for a single $\mathsf{W}$-manifold: proof
Let $d\in\mathbb{Z}^{\geq 1}$, $y\in\mathbb{R}$ and $A\in\mathbb{R}^{>0}$. Define the functions $N,M:\mathbb{R}\longrightarrow\mathbb{R}$ as
$N(\lambda)$ $:=\sum_{\begin{subarray}{c}m-y\in\mathbb{Z}^{\geq 0}\\ Am^{2}\leq\lambda^{2}\end{subarray}}m^{d-1}\qquad\textnormal{and}\qquad M(\lambda):=\frac{1}{d(\sqrt{A})^{d}}\lambda^{d}+\frac{1}{2(\sqrt{A})^{d-1}}\lambda^{d-1}.$
Then $N\sim_{2}M$.
Proof.
Let us denote $\tilde{m}:=m-y$. Then $\tilde{m}\in\mathbb{Z}^{\geq 0}$ and the condition $Am^{2}\leq\lambda^{2}\Leftrightarrow-\frac{\lambda}{\sqrt{A}}-y\leq\tilde{m}\leq\frac{\lambda}{\sqrt{A}}-y$, owing to the fact $\tilde{m}\in\mathbb{Z}^{\geq 0}$, becomes $0\leq\tilde{m}\leq\left[\frac{\lambda}{\sqrt{A}}-y\right]$.
Let $\gamma(\lambda):=\frac{\lambda}{\sqrt{A}}-y$. Then
$N(\lambda)$ $=\sum_{\tilde{m}=0}^{[\gamma(\lambda)]}(\tilde{m}+y)^{d-1}$ (62)
$=\sum_{\tilde{m}=0}^{[\gamma(\lambda)]}\left(\tilde{m}^{d-1}+(d-1)\tilde{m}^{d-2}y+\cdots+y^{d-1}\right).$
Using corollary 7.2, we note that
$\sum_{\tilde{m}=0}^{[\gamma(\lambda)]}\tilde{m}^{d-1}=\frac{1}{d}[\gamma(\lambda)]^{d}+\frac{1}{2}[\gamma(\lambda)]^{d-1}+\text{terms with lower powersof $[\gamma(\lambda)]$},$
and
$\sum_{\tilde{m}=0}^{[\gamma(\lambda)]}(d-1)\tilde{m}^{d-2}y=y[\gamma(\lambda)]^{d-1}+\frac{1}{2}(d-1)y[\gamma(\lambda)]^{d-2}+\text{terms with lower powers of $[\gamma(\lambda)]$}.$
Using this in (62), we have
$N(\lambda)$ $=\widetilde{N}(\gamma(\lambda)),\qquad\textnormal{where}$
$\widetilde{N}(\gamma)$ $=\frac{1}{d}[\gamma]^{d}+\left(\frac{1}{2}+y\right)[\gamma]^{d-1}+\text{terms with lower powers of $[\gamma]$}.$
Let us now define $\widetilde{M}:\mathbb{R}\longrightarrow\mathbb{R}$ as
$\widetilde{M}(\gamma)$ $:=\frac{1}{d}\gamma^{d}+\left(\frac{1}{2}+y\right)\gamma^{d-1}.$
By proposition 7.5, $\widetilde{N}\sim_{2}\widetilde{M}$. As a result, using proposition 7.6, we conclude that $N\sim_{2}M$, whence the claim is proved. ∎
Remark 7.8.
Note that in the expression for $M(\lambda)$, the coefficient of $\lambda^{d-1}$ is non-zero. Hence, theorem 7.7 shows that the error term for $N(\lambda)$ can not be improved beyond $\lambda^{d-1}$ in addition to being no worse than $\lambda^{d-1}$.
Using the above theorem, we next prove the following result.
Theorem 7.9.
Let $d\in\mathbb{Z}^{\geq 2}$ and $\mathsf{Q}_{d-1}(t)$ be a degree $d-1$ polynomial in $t$ that does not have the $t^{d-2}$ term. In other words,
@@
where $\varphi(t)$ is a polynomial of degree at most $d-3$ and $C$ is some non-zero constant. Let $y$ be a fixed real number and $A$ a fixed positive real numbers. Define $N,M:\mathbb{R}\longrightarrow\mathbb{R}$ as
$N(\lambda)$ $:=\sum_{\begin{subarray}{c}m-y\in\mathbb{Z}^{\geq 0}\\ Am^{2}\leq a\lambda^{2}+b\end{subarray}}\mathsf{Q}_{d-1}(m)\qquad\textnormal{and}$
$M(\lambda)$ $:=\frac{Ca^{\frac{d}{2}}}{d\left(\sqrt{A}\right)^{d}}\lambda^{d}+\frac{Ca^{\frac{d-1}{2}}}{2\left(\sqrt{A}\right)^{d-1}}\lambda^{d-1},$
where $a\in\mathbb{R}^{>0}$ and $b\in\mathbb{R}$. Then $N\sim_{2}M$.
@@
This follows from theorem 7.7. By applying binomial theorem to $\Big{(}\frac{a\lambda^{2}+b}{A}\Big{)}^{\frac{d}{2}}$, the leading term gives us the desired asymptotics. The lower order terms can not cancel off or change the error term, because the $t^{d-2}$ term is absent in $\mathsf{Q}_{d-1}(t)$. ∎∎
Proof of theorem 3.3: Follows immediately from theorem 7.9. ∎
@@ 8. Acknowledgement
We are grateful to Antareep Saud for several useful discussions related to this project. The third author is also grateful to Vamsi Pingali for helpful discussions and comments and in particular pointing out the reference [2] for the proof of Weyl’s law.

### v2

@@ 7. Error is sharp for a single $\mathsf{W}$-manifold: proof
Let $d\in\mathbb{Z}^{\geq 2}$ and $\mathsf{Q}_{d-1}(t)$ be a degree d1d-1 polynomial in $t$ that does not have the $t^{d-2}$ term. In other words,
@@
where $\varphi(t)$ is a polynomial of degree at most d3d-3 and $C$ is some non-zero constant. Let $y$ be a fixed non negative real number and $A$ a fixed positive real numbers. Define $N:\mathbb{R}\longrightarrow\mathbb{R}$ as
$N(\lambda)$ $:=\sum_{\begin{subarray}{c}m-y\in\mathbb{Z}^{\geq 0}\\ Am^{2}\leq a\lambda^{2}+b\end{subarray}}\mathsf{Q}_{d-1}(m).$ (86)
where $a\in\mathbb{R}^{>0}$ and $b\in\mathbb{R}$. Then the error term in the asymptotic expression of $N(\lambda)$ (after subtracting off the leading term) can not be improved by $O(\lambda^{d-1-\varepsilon})$ for any positive $\varepsilon$.
@@
Without loss of generality, let us set $C$ to be equal to one. We sum up the leading term of $\mathsf{Q}_{d-1}(t)$ (namely the degree $t^{d-1}$ terms) using Proposition 7.1. The first two terms precisely correspond to $M(\lambda)$ as given by Lemma 7.2 (replace $a$ by $\frac{a}{A}$ and replace $b$ by $\frac{b}{A}$). The remaining terms are all of order $\lambda^{d-2}$ or less (this is where we use the fact that $\mathsf{Q}_{d-1}(t)$ does not have a $t^{d-2}$ term). Hence the remaining terms can not compensate for the error term from $M(\lambda)$. We have already seen that the error term for $M(\lambda)$ can not be improved. This shows that the error term for $N(\lambda)$ can not be improved. ∎
Since the Eigenvalue counting function $N(\lambda)$ for all $\mathsf{W}$-manifolds is of the form considered in Theorem 7.3, we have proved Theorem 1.3
8. Counting Lattice Points in a Ball
To keep the paper self contained, we give a proof of equation (27) (even though it is a standard fact). We will prove the following general result.
Proposition 8.1.
Let $p$ be any norm on $\mathbb{R}^{d}$ where $d\geq 1$. Given $\lambda>0$, let $B_{p}(\lambda)$ denote the closed $p$-ball of radius $\lambda$ around the origin, i.e. $B_{p}(\lambda):=\{x\in\mathbb{R}^{d}:p(x)\leq\lambda\}$. Let $N^{p}(\lambda):=\#(\mathbb{Z}^{d}\cap B_{p}(\lambda))$ denote the number of integer lattice points in $B_{p}(\lambda)$. Then, $N^{p}(\lambda)\in O(\lambda^{d})$ as $\lambda\to\infty$.
The supremum norm on $\mathbb{R}^{d}$ is particularly suited for counting integer lattice points, so this is the case we first deal with.
Lemma 8.2.
Let $p_{\infty}$ denote the supremum norm on $\mathbb{R}^{d}$, i.e. $p_{\infty}(x):=\max_{1\leq i\leq n}|x_{i}|$. Then $\forall\lambda>0$, $N^{p_{\infty}}(\lambda)\leq(2\lambda+3)^{d}$.
We will simplify notation slightly: $B_{\infty}(\lambda)$ will denote $B_{p_{\infty}}(\lambda)$ and $N^{\infty}(\lambda)$ will denote $N^{p_{\infty}}(\lambda)$.
Proof.
Observe that $B_{\infty}(\lambda)=[-\lambda,\lambda]^{d}$ is a $d$-dimensional cube centered at the origin.
We will first assume $\lambda\in\mathbb{N}$. It is then clear that
$x\in\mathbb{Z}^{d}\cap B_{\infty}(\lambda)\iff x_{i}\in\mathbb{Z}\cap[-\lambda,\lambda]\text{ for each }1\leq i\leq d$
As $\lambda\in\mathbb{N}$, this means each $x_{i}$ can take one of precisely $2\lambda+1$ values, i.e
$x_{i}\in\{-\lambda,-\lambda+1,\ldots,0,\ldots,\lambda-1,\lambda\}$
and we conclude
$N^{\infty}(\lambda)=(2\lambda+1)^{d}\text{ whenever }\lambda\in\mathbb{N}$ (87)
Now, let $\lambda>1$ be any real number. Letting $\lfloor\lambda\rfloor$ denote the greatest integer less than or equal to $\lambda$, it is clear that
$N^{\infty}(\lambda)=N^{\infty}(\lfloor\lambda\rfloor)\leq N^{\infty}(\lfloor\lambda\rfloor+1)=(2\lfloor\lambda\rfloor+3)^{d}$
where the last equality follows from (87) with $\lfloor\lambda\rfloor+1$ instead of $\lambda\in\mathbb{N}$.
Finally, we have
$N^{\infty}(\lambda)\leq(2\lfloor\lambda\rfloor+3)^{d}\leq(2\lambda+3)^{d}$
where the second inequality follows from the fact that $\lfloor\lambda\rfloor\leq\lambda$ and $x\to(2x+3)^{d}$ is an increasing function on $\mathbb{R}_{+}$. ∎
With Lemma 8.2 in hand, the proof of Proposition 8.1 follows easily.
Proof of Proposition 8.1: As all norms are equivalent on $\mathbb{R}^{d}$, we have positive constants $m,M$ such that for any $x\in\mathbb{R}^{d}$
$mp(x)\leq p_{\infty}(x)\leq Mp(x)$ (88)
We make a simple observation: for any $\lambda>0$
$B_{p}(\lambda)\subset B_{\infty}(M\lambda)$ (89)
This follows immediately from the second inequality in (88).
From (89) and the definitions of $N^{p}$ and $N^{\infty}$ we get that for any $\lambda>0$,
$N^{p}(\lambda)\leq N^{\infty}(M\lambda)$ (90)
Applying Lemma 8.2 to $N^{\infty}(M\lambda)$, we get that for any $\lambda>0$
$N^{\infty}(M\lambda)\leq(2M\lambda+3)^{d}$ (91)
Finally, (90) and (91) combine to give that for any $\lambda>0$,
$N^{p}(\lambda)\leq N^{\infty}(M\lambda)\leq(2M\lambda+3)^{d}$
And that is enough to show that $N^{p}(\lambda)\in O(\lambda^{d})$ as $\lambda\to\infty$
@@ 9. Acknowledgement
We are grateful to Antareep Saud for several useful discussions related to this project. The second author is also grateful to Vamsi Pingali for helpful discussions and comments and in particular pointing out the reference [2] for the proof of Weyl’s law.